# QCD Phase Diagram and the Finite Volume Fireball: A Model Study

### v2

@@ I Introduction
In this paper, the aim is to get an idea of the effect of the finite volume on the phase diagram for a static system; the effects of the dynamic nature is being left for a future study. We study the phase diagram using the Nambu Jona-Lasinio (NJL) model [10], which is an effective model expected to capture the physics related to the chiral symmetry, in particular, the spontaneous chiral symmetry breaking. Since the finite temperature crossover is associated with the melting of the chiral condensate, the NJL model is expected to capture the basic features of the crossover.
@@ II Formalism and the boundary condition
The four-fermion term in $\mathcal{L}_{\scriptscriptstyle\rm NJL}$ makes the theory non-renormalizable. Hence, a suitable regularization method has to be specified, and the results will depend on the regularization parameter. We have used Schwinger’s proper time regularization [27, 28], with the cutoff $\tau_{\rm\scriptscriptstyle UV}=1/\Lambda_{\rm\scriptscriptstyle UV}^{2}$. Therefore, our model has three parameters: $m,G$ and the UV momentum cutoff $\Lambda_{\rm\scriptscriptstyle UV}$. These parameters can be chosen by specifying the pion mass, the pion decay constant and either the chiral condensate or the chiral transition temperature ($T_{\chi}$). Some suitable parameter sets, and how they affect the chiral crossover temperature, have been discussed in Ref. [28]. The phase diagram can be qualitatively different for the different parameter sets. In particular, for some choices of parameters one gets a critical point and a first-order transition at large quark number chemical potential $\mu=\dfrac{\mu_{\scriptscriptstyle B}}{3}$, while for some other choices, the transition is a crossover in the whole $\left(T,\,\mu\right)$ plane. We have chosen the following parameter set:
@@
where $\hat{n}$ is the unit normal four-vector to the boundary. The allowed momentum modes for Eq. (5) are well-known; the relevant equations are given in Appendix A. Then the expression for $\Omega$ becomes:
$\Omega=\frac{(M-m)^{2}}{4\,G}-\frac{N_{c}\,N_{f}}{V}\sum_{j}\sum_{p_{j_{\kappa}}}(2j+1)\left[\left\{-\frac{\Lambda\,e^{-\left(\frac{E}{\Lambda}\right)^{2}}}{\sqrt{\pi}}+E\,\text{Erfc}\left(\frac{E}{\Lambda}\right)\right\}+T\,\text{log}\left(1+e^{-\frac{(E-\mu_{q})}{T}}\right)+T\,\text{log}\left(1+e^{-\frac{(E+\mu_{q})}{T}}\right)\right],$ (7)
@@
The finite temperature phase transition is chiral symmetry restoring in nature, and we will track the transition using $\langle\bar{\psi}\psi\rangle$, or equivalently, $M$. In the absence of the symmetry-breaking term $m$, in the full quantum theory, the chiral symmetry will be restored in finite volume at all temperatures, due to tunnelling between the classical vacua of $\mathcal{L}^{\scriptscriptstyle\rm MF}_{\scriptscriptstyle\rm NJL}$. In Monte Carlo simulations, this will show an instability in $\langle\bar{\psi}\psi\rangle$. A mean-field treatment will miss the physics of tunnelling, and indicate symmetry breaking at low temperatures. This problem is not there in the presence of the mass term corresponding to physical light quarks, where lattice simulations show a stable result for $\langle\bar{\psi}\psi\rangle$ at quite small volumes down to $m_{\pi}L\sim 1$.
@@ III Chiral transition at finite volume
Now we introduce a nonzero quark chemical potential $\mu$. For small values of $\mu$, the qualitative behavior is very similar to what is seen for $\mu=0$. In Fig. 2, we show $M$ as a function of $T$ at different volumes for three values of $\mu$. The temperature and volume dependence of $M$ at $\mu$=100 MeV is very similar to that at $\mu$=0, as a comparison of the left panel of Fig. 2 with Fig. 1 shows. As $\mu$ increases, the crossover shifts to smaller values of temperature, as expected. Also, we find a stiffening of the transition at larger volumes: the smoothing of the transition at small volumes is enhanced. At $\mu$=350 MeV, shown in the right panel of Fig. 2, we find a discontinuity, indicating a first-order transition, at infinite volume. The first-order nature of the transition can be further ascertained by looking at $\Omega$, Fig. 3. At the transition region, a two-minimum structure is seen, which is the hallmark of first-order transition, and indicates phase coexistence. At this $\mu$ the transitions are smooth in the finite volume system, even at $R$=20 fm, though at such a large volume it is quite a sharp crossover. We also find here a much more enhanced smoothing of the transition with a reduction of volume: while the curve at $R$ = 20 fm indicates a sharp crossover, the one at $R$=5 fm indicates a smooth behavior. Also, the crossover temperature changes much more rapidly here with volume: the transition temperature changes from 47.3 MeV in the infinite volume system to 72.2 MeV in the 5 fm sphere.
@@
Figure 2: The dependence of the $M$ vs $T$ curve with volume. Results at $\mu$=100 MeV (left), 250 MeV (middle) and 350 MeV (right) is shown. A discontinuous transition is seen in the infinite volume system at $\mu$=350 MeV.
@@
Before proceeding with the discussion of the investigation of the transition line at higher $\mu$ at finite volume, let us clarify our terminology. In our analysis, the term “first-order transition” corresponds to a two-well structure in the free energy, with well-separated local minima at small and large values of $M$ corresponding to the symmetry restored and symmetry broken phases, respectively. Such a structure will indicate phase coexistence, and possible supercooling/superheating in a dynamical system. This is what we have looked for, and have referred to as the signature of a first-order transition.
In the infinite volume case, tunnelling between the two minima is prohibited. However, in a finite volume, the free energy barrier between the minima is finite, and there is a nonzero probability for tunnelling. This will lead to a unique minimum at every $\left(T,\,\mu\right)$ value for a static system, and strictly speaking, one can only have crossovers in the finite volume system. However, the tunnelling probability is exponentially suppressed in volume. In a fireball which rapidly cools through the transition region, tunnelling will be insignificant unless the volume is very small. In cases where the classical free energy shows a two-minima structure and a discreet jump at the transition, we will expect interesting phenomena, related to the features of a first-order transition, at the transition point (see also Ref.[15]). Therefore, the two-well structure of the mean-field potential, $\Omega$, is of actual physical interest, and we will loosely use terms like first-order line, phase coexistence etc. to refer to such a structure even in a finite volume system. An analysis of the dynamics at discontinuity in a finite volume system has been done in Ref. [32].
@@
We illustrate this behavior in Fig. 4, where we show $\Omega$ at different volumes at $T$=20 MeV. We find a two-well structure in $\Omega$ even at $R$=10 fm. Note also the range of $\mu$ over which this structure is seen (we loosely refer below to this region as the coexistence region). A shift of the transition to higher $\mu$ is seen with the reduction in volume; at the same time, a weakening of the double-well structure (and a shrinking of the coexistence region) is seen. At this temperature, we do not find a discontinuous transition in the 5 fm sphere, but as we go further down to $T$=0 we find a discontinuous transition in all our systems.
@@ IV Number density and quark number susceptibility
In the case of a conventional boundary condition like the antiperiodic one, the shift of the endpoint of first order line is insignificant for a box of size 5 fm or higher (see Appendix B), and the change in the magnitude of the $\chi_{2}$ peaks are due to the smoother nature of the transition at smaller volumes. As Fig. 8 and Fig. 9 show, here the major effect on the peak structure is due to the shift of the transition point. While our calculation is only at the mean field level, we believe this effect needs to be kept in mind when looking at the volume dependence of the baryon number cumulants (which are closely related to the susceptibilities).
@@ Appendix A Calculational details
$\Omega=\frac{(M-m)^{2}}{4\,G}-2\,N_{c}\,N_{f}\int\dfrac{d^{3}p}{(2\pi)^{3}}\left[\left\{-\frac{\Lambda\,e^{-\left(\frac{E}{\Lambda}\right)^{2}}}{\sqrt{\pi}}+E\,\text{Erfc}\left(\frac{E}{\Lambda}\right)\right\}+T\,\text{log}\left(1+e^{-\frac{(E-\mu_{q})}{T}}\right)+T\,\text{log}\left(1+e^{-\frac{(E+\mu_{q})}{T}}\right)\right].$ (15)
@@
where $q_{\kappa,i}\equiv pR$ is obtained by solving Eq. (17). Here, $i$ is the radial excitation quantum number ($i=1,2,3,...$) and $\kappa=\pm(j+1/2)$ is for the total angular momentum, $j=1/2,3/2,...$. Replacing the three-momentum integral in Eq. (15) by a sum over these modes leads to Eq. (7).
@@ Appendix B Comparison with other boundary conditions
In the right panel of the figure, we have compared the phase transition line in the $\left(T,\,\mu\right)$ plane for an infinite volume system, along with those obtained for a finite volume box of dimension 5 fm with APBC and a sphere of radius 5 fm with MIT boundary condition. The interpretation of the line is similar to that used in Section III and in Fig. 5: it is the line obtained from the $M$ dependence of the mean field effective potential. The antiperiodic box shows very little volume dependence even at 5 fm: the crossover line is very close to the infinite volume one, and even the coexistence region in $M$ starts to appear at a very similar values of $\left(T,\,\mu\right)$  [^2] For the linear sigma model, a significant shift of the transition line, and shift of the endpoint of the first order line towards lower temperatures, have been reported for a 5 fm box with APBC and PBC [15].. At lower temperatures, the “first order line” (in the mean field effective potential sense) shows slight deviation from the infinite volume case, but still within the coexistence region. On the other hand, as discussed in Section III, the MIT boundary condition leads to a substantial change: the crossover line is pushed to higher values of $\mu$ at all temperatures. Also the coexistence region is shrunk significantly at this volume. Fig. 10 supports our intuition that the studies with the APBC and similar boundary conditions significantly underestimate the finite volume effect.

### v3

@@ I Introduction
In this paper, the aim is to get an idea of the effect of the finite volume on the phase diagram for a static system; the effects of the dynamic nature are being left for a future study. We study the phase diagram using the Nambu Jona-Lasinio (NJL) model [10], which is an effective model expected to capture the physics related to the chiral symmetry, in particular, the spontaneous chiral symmetry breaking. Since the finite temperature crossover is associated with the melting of the chiral condensate, the NJL model is expected to capture the basic features of the crossover.
@@ II Formalism and the boundary condition
The four-fermion term in $\mathcal{L}_{\scriptscriptstyle\rm NJL}$ makes the theory non-renormalizable. Hence, a suitable regularization method has to be specified, and the results will depend on the regularization parameter. We have used Schwinger’s proper time regularization [27, 28], with the cutoff $\tau_{\rm\scriptscriptstyle UV}=1/\Lambda_{\rm\scriptscriptstyle UV}^{2}$. Therefore, our model has three parameters: $m,G$ and the UV momentum cutoff $\Lambda_{\rm\scriptscriptstyle UV}\equiv\Lambda$. These parameters can be chosen by specifying the pion mass, the pion decay constant and either the chiral condensate or the chiral transition temperature ($T_{\chi}$). Some suitable parameter sets, and how they affect the chiral crossover temperature, have been discussed in Ref. [28]. The phase diagram can be qualitatively different for the different parameter sets. In particular, for some choices of parameters one gets a critical point and a first-order transition at large quark number chemical potential $\mu=\dfrac{\mu_{\scriptscriptstyle B}}{3}$, while for some other choices, the transition is a crossover in the whole $\left(T,\,\mu\right)$ plane. We have chosen the following parameter set:
@@
where $\hat{n}$ is the unit normal four-vector to the boundary. The allowed momentum modes for Eq. (5) are well-known; the relevant equations are given in Appendix A. Then the expression for the mean-field thermodynamic potential $\Omega$ becomes:
$\Omega=\frac{(M-m)^{2}}{4\,G}-\frac{N_{c}\,N_{f}}{V}\sum_{j}\sum_{p_{j_{\kappa}}}(2j+1)\left[\left\{-\frac{\Lambda\,e^{-\left(\frac{E}{\Lambda}\right)^{2}}}{\sqrt{\pi}}+E\,\text{Erfc}\left(\frac{E}{\Lambda}\right)\right\}+T\,\text{log}\left(1+e^{-\frac{(E-\mu)}{T}}\right)+T\,\text{log}\left(1+e^{-\frac{(E+\mu)}{T}}\right)\right],$ (7)
@@
The finite temperature phase transition is chiral symmetry restoring in nature, and we will track the transition using $\langle\bar{\psi}\psi\rangle$, or equivalently, $M$. In the absence of the symmetry-breaking term $m$, in the full quantum theory, the chiral symmetry will be restored in finite volume at all temperatures, due to tunneling between the classical vacua of $\mathcal{L}^{\scriptscriptstyle\rm MF}_{\scriptscriptstyle\rm NJL}$. In Monte Carlo simulations, this will show an instability in $\langle\bar{\psi}\psi\rangle$. A mean-field treatment will miss the physics of tunneling, and indicate symmetry breaking at low temperatures. This problem is not there in the presence of the mass term corresponding to physical light quarks, where lattice simulations show a stable result for $\langle\bar{\psi}\psi\rangle$ at quite small volumes down to $m_{\pi}L\sim 1$.
@@ III Chiral transition at finite volume
Now we introduce a nonzero quark chemical potential $\mu$. For small values of $\mu$, the qualitative behavior is very similar to what is seen for $\mu=0$. In Fig. 2, we show $M$ as a function of $T$ at different volumes for three values of $\mu$. The temperature and volume dependence of $M$ at $\mu$=100 MeV is very similar to that at $\mu$=0, as a comparison of the left panel of Fig. 2 with Fig. 1 shows. As $\mu$ increases, the crossover shifts to smaller values of temperature, as expected. Also, we find a stiffening of the transition at larger volumes: the smoothing of the transition at small volumes is enhanced. At $\mu$=350 MeV, shown in the right panel of Fig. 2, we find a discontinuity, indicating a first-order transition, at infinite volume. The first-order nature of the transition can be further ascertained by looking at $\Omega$, Fig. 3. At the transition region, a two-minima structure is seen, which is the hallmark of first-order transition, and indicates phase coexistence. At this $\mu$ the transitions are smooth in the finite volume system, even at $R$=20 fm, though at such a large volume it is quite a sharp crossover. We also find here a much more enhanced smoothing of the transition with a reduction of volume: while the curve at $R$ = 20 fm indicates a sharp crossover, the one at $R$=5 fm indicates a smooth behavior. Also, the crossover temperature changes much more rapidly here with volume: the transition temperature changes from 47.3 MeV in the infinite volume system to 72.2 MeV in the 5 fm sphere.
@@
Figure 2: The dependence of the $M$ vs $T$ curve with volume. Results at $\mu$=100 MeV (left), 250 MeV (middle) and 350 MeV (right) are shown. A discontinuous transition is seen in the infinite volume system at $\mu$=350 MeV.
@@
Before proceeding with the discussion of the investigation of the transition line at higher $\mu$ at finite volume, let us clarify our terminology. In our analysis, the term “first-order transition” corresponds to a two-well structure in the mean-field free energy, with well-separated local minima at small and large values of $M$, corresponding to the symmetry restored and symmetry broken phases, respectively. At the transition point, e.g., $(T,\mu)\approx(47.3,350)$ MeV for the infinite volume case shown in Fig. 3, $M$ shows a discontinuous jump. Such a two-minima structure is associated with phase coexistence, and possible supercooling/superheating in a dynamical system. This is what we have looked for, and have referred to as the signature of a first-order transition. Similar structures at finite volume are seen at larger $\mu$ (see Fig. 4).
Beyond mean-field level, we have to include fluctuations. In particular, in a finite volume system, the free energy barrier between the two local minima is finite, and there is a nonzero probability for tunneling $\sim\exp(-V\Delta\Omega/T)$ between them. This will smoothen the discontinuity in a static system. However, the tunneling rate is exponentially suppressed in volume. In a fireball which rapidly cools through the transition region, tunneling will be insignificant unless the volume is very small. In cases where the mean field free energy shows a two-minima structure and a discrete jump at the transition, we will expect interesting phenomena, related to the features of a first-order transition, at the transition point (see also Ref.[15]). Therefore, the two-well structure of the mean-field potential, $\Omega$, is of actual physical interest, and we will loosely use terms like first-order line, phase coexistence etc. to refer to such a structure even in a finite volume system. An analysis of the dynamics at discontinuity in a finite volume system has been done in Ref. [32].
@@
In Fig. 4, $\Omega$ is shown at $T$=20 MeV for different volumes. We find a two-well structure in $\Omega$ even at $R$=10 fm. Note also the range of $\mu$ over which this structure is seen (we loosely refer below to this region as the coexistence region). A shift of the transition to higher $\mu$ is seen with the reduction in volume; at the same time, a weakening of the double-well structure (and a shrinking of the coexistence region) is seen. At this temperature, we do not find a discontinuous transition in the 5 fm sphere, but as we go further down to $T$=0 we find a discontinuous transition in all our systems.
@@ IV Number density and quark number susceptibility
In the case of a conventional boundary condition like the antiperiodic one, the shift of the endpoint of first order line is insignificant for a box of size 5 fm or higher (see Appendix B), and the change in the magnitude of the $\chi_{2}$ peaks is due to the smoother nature of the transition at smaller volumes. As Fig. 8 and Fig. 9 show, here the major effect on the peak structure is due to the shift of the transition point. While our calculation is only at the mean field level, we believe this effect needs to be kept in mind when looking at the volume dependence of the baryon number cumulants (which are closely related to the susceptibilities).
@@ Appendix A Calculational details
$\Omega=\frac{(M-m)^{2}}{4\,G}-2\,N_{c}\,N_{f}\int\dfrac{d^{3}p}{(2\pi)^{3}}\left[\left\{-\frac{\Lambda\,e^{-\left(\frac{E}{\Lambda}\right)^{2}}}{\sqrt{\pi}}+E\,\text{Erfc}\left(\frac{E}{\Lambda}\right)\right\}+T\,\text{log}\left(1+e^{-\frac{(E-\mu)}{T}}\right)+T\,\text{log}\left(1+e^{-\frac{(E+\mu)}{T}}\right)\right].$ (15)
@@
where $q_{\kappa,i}\equiv pR$ is obtained by solving Eq. (17). Here, $i$ is the radial excitation quantum number ($i=1,2,3,...$) and $\kappa=\pm(j+1/2)$, where $j=1/2,3/2,...$ is the total angular momentum quantum number. Replacing the three-momentum integral in Eq. (15) by a sum over these modes leads to Eq. (7).
@@ Appendix B Comparison with other boundary conditions
In the right panel of the figure, we have compared the phase transition line in the $\left(T,\,\mu\right)$ plane for an infinite volume system, along with those obtained for a finite volume box of dimension 5 fm with APBC and a sphere of radius 5 fm with MIT boundary condition. The interpretation of the line is similar to that used in Section III and in Fig. 5: it is the line obtained from the $M$ dependence of the mean field effective potential. The antiperiodic box shows very little volume dependence even at 5 fm: the crossover line is very close to the infinite volume one, and even the coexistence region in $M$ starts to appear at a very similar value of $\left(T,\,\mu\right)$  [^2] For the linear sigma model, a significant shift of the transition line, and shift of the endpoint of the first order line towards lower temperatures, have been reported for a 5 fm box with APBC and PBC [15].. At lower temperatures, the “first order line” (in the mean field effective potential sense) shows slight deviation from the infinite volume case, but still within the coexistence region. On the other hand, as discussed in Section III, the MIT boundary condition leads to a substantial change: the crossover line is pushed to higher values of $\mu$ at all temperatures. Also the coexistence region is shrunk significantly at this volume. Fig. 10 supports our intuition that the studies with the APBC and similar boundary conditions significantly underestimate the finite volume effect.